HUBBLE SPACE TELESCOPE DAILY REPORT #2410 PERIOD COVERED: 0000Z (UTC) 07/08/99 - 0000Z (UTC) 07/09/99 Daily Status Report as of 190/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8444 (WF/PC-2 CYCLE-8 Internal Monitor) The WF/PC-2 was used to make a bias, and an internal flat, observation as part of a Cycle-8 monitoring program to monitor the health of the cameras. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.3 Completed Two Sets of WF/PC-2/STIS/CCD 8267 (Taking the Measure of Planets in the Globular Cluster 47 Tucanae) Both the WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to make observations of the globular cluster NGC-104 (a.k.a., Globular Cluster 47 Tucanae) to search for stars with planetary companions in close orbits. HST is uniquely capable of providing the requisite, high-dynamic range, time-series photometry. These observations extended over a several day period. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 18 Successful : 18 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Operations Notes: The flight operation team returned to the STOCC area at 190/0000Z following the conclusion of GSFC Internal Simulation #7 (EVA-2). They continue to operate using the PRS ground system. The “Tucanae” proposal continues to be executed with no reported problems. Using ROP SR-1, the SSR EDAC error counter was cleared four times. The NICMOS FPA temperature sensor for detector 2 was enabled at 189/1745Z as directed by an operations request. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH